OpenStudy (anonymous):

A box shaped like a rectangular prism has a height of 17 in. And a volume of 2720 in^3.the length is 4 inches greater than twice the width. What is the width of the box?

OpenStudy (anonymous):

Please help!!! I need a good grade!

OpenStudy (anonymous):

So we have a height of 17 in., a length that is 2w+4 in., and a width of w in. V=(17)(2w+4)(w) We know V=2720 2720=(17)(2w+4)(w) Then just solve for w

OpenStudy (anonymous):

Thnk you :-)